OpenStudy (anonymous):

convert from polar to rectangular r=4/(5-3 sin (theta)) I got (25/16 X^2) + y^2 - (24/16 y) = 1 then i am stuck need to be in standard form of ellipse ( im guessing ellipse but could not get the form right) or hyperbola

zepdrix (zepdrix):

Should work out to an ellipse? Mmm ok let's see if we can figure this out...

OpenStudy (anonymous):

I can see that it is an ellipse but i cant get it to be like the standard from as i need to graph it

zepdrix (zepdrix):

\[\Large r\quad=\quad \frac{4}{5-3\sin \theta}\]So doing some multiplication:\[\Large 5r-3r \sin \theta\quad=\quad 4\]We can covert the r and r sin theta into some stuff,\[\Large 5\sqrt{x^2+y^2}-3y\quad=\quad 4\]Make sense so far? :O Look ok?

OpenStudy (anonymous):

yup exactly what i got

zepdrix (zepdrix):

Hmmm let's see..\[\Large 5\sqrt{x^2+y^2}\quad=\quad 4+3y\]Squaring:\[\Large 5x^2+5y^2\quad=\quad 9y^2+24y+16\]Then we have to simplify this down? :O Mmm doesn't look too bad :D

OpenStudy (anonymous):

I got (25/16 X^2) + y^2 - (24/16 y) = 1 then i am stuck

OpenStudy (anonymous):

5x2+5y2 shouldn't the 5 be squared as well since we are suppose to square everything right?

zepdrix (zepdrix):

Ah yes, good catch! ^^

zepdrix (zepdrix):

Mmm ok lemme fix that a sec before I write anything else lol.\[\Large 25x^2+25y^2\quad=\quad 9y^2+24y+16\]

OpenStudy (anonymous):

alright i got that too...

zepdrix (zepdrix):

Ya lemme erase that garbage a sec :x let's do that a smarter way

zepdrix (zepdrix):

Moving the y's to the left gives us: \[\Large 25x^2+\color{#CC0033}{16y^2-24y}\quad=\quad16\] Now we need to complete the square on the y's before we do anything else.

OpenStudy (anonymous):

ok im doing it too hope it come out to the standard form

OpenStudy (anonymous):

yeah i am stuck \[25x ^{2}+16(y+\frac{ 3 }{ 4 })^{2}=16+\frac{ 9 }{ 16 }\]

zepdrix (zepdrix):

\[\Large 25x^2+16\left(\color{#CC0033}{y^2-\frac{3}{2}y}\right)\quad=\quad16\] \[\Large 25x^2+16\left(\color{#CC0033}{y^2-\frac{3}{2}y+\frac{9}{16}}\right)\quad=\quad16\color{#CC0033}{+9}\] \[\Large 25x^2+16\left(\color{#CC0033}{y-\frac{3}{4}}\right)^2\quad=\quad25\]

zepdrix (zepdrix):

So when you added 9/16 to complete the square, you should also subtract 9/16 `within the brackets`. So to pull that value out, you need to multiply it by 16, right? Because of the 16 on the brackets?

OpenStudy (anonymous):

oh yeah....

zepdrix (zepdrix):

So I think then we just need to divide by 25, yes? :)

zepdrix (zepdrix):

We might have to do a little work with the y term so it looks more "standard".

OpenStudy (anonymous):

How do we get the x term to be \[(x-h)^{2}\]

zepdrix (zepdrix):

\[\Large x^2+\frac{16\left(\color{#CC0033}{y-\frac{3}{4}}\right)^2}{25}\quad=\quad1\] \[\Large x^2+\frac{\left(y-\frac{3}{4}\right)^2}{\left(\frac{5}{4}\right)^2}\quad=\quad1\]So like that for the y's maybe?

zepdrix (zepdrix):

The x is already in standard form:\[\Large \frac{(x-0)^2}{1^2}\quad=\quad x^2\]See? :o

OpenStudy (anonymous):

Okay now i see

zepdrix (zepdrix):

Hmm I hope we did that correctly :) Bit of a tricky problem to simplify.

OpenStudy (anonymous):

i checked everything as we progress should be fine thank you so much

zepdrix (zepdrix):

cool c: